OpenStudy (anonymous):

What is the solution set for the open sentence with the given replacement set? 4y + 2 = 18, {4, 5, 6, 7} A. 4 B. 5 C. 6 D. 7

pooja195 (pooja195):

4(4) + 2 = 18 4(5) + 2 = 18 4(6) + 2 = 18 4(7) + 2 = 18 Which one makes a true statement?

OpenStudy (anonymous):

when i do it i get a high number

pooja195 (pooja195):

lets do it in order... 4(4) + 2 = 18 Whats 4 x 4 ?

OpenStudy (anonymous):

16

pooja195 (pooja195):

16 + 2 = ?

OpenStudy (anonymous):

18

pooja195 (pooja195):

Congrats you just found the answer :)

OpenStudy (anonymous):

and this answer was a right?

pooja195 (pooja195):

Yep

OpenStudy (anonymous):

thx

pooja195 (pooja195):

:) \(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e\color{#24ff00}~\color{#00ff00}t\color{#00ff49}o\color{#00ff92}~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}S\color{#0000ff}t\color{#2400ff}u\color{#6D00ff}d\color{#B600ff}y\color{#FF00ff}!\)\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Large\bf\color{Lime}{Please,\ don't\ hesitate\ to\ ask\ questions!}\)\(\Large\bf\color{magenta}{We're\ here\ to\ help!}\Large\color{magenta}{\ddot\smile}\)